OpenStudy (anonymous):

please help me solve this question a 55g sample of gaseous fuel mixture contains 0.51mole fraction of propane C3H8 ,the remainder of the mixture is butane C4H10 . what are the masses for propane and butane in the sample ?

OpenStudy (mos1635):

set n =total moles of mixture n1 = 0.51*n moles of propane n2=0.49*n moles of butane M(total)=55 M1+M2=55 n1*MM1+n2*MM2=55 substitude solve to n calc M1 and M2...